Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for October 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, October 28, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 21, 1998 to 7:00 a.m. Wednesday, Oct. 28, 1998.

The Thermal Destruction Facility (TDF) was on-line for 166.3 hours and processed 6565.8 tons of contaminated soil during the time frame of this report. The TDF was on-line 98.4 % of the time with an average hourly feed rate of 39.5 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 167,216 tons of contaminated soil, which amounts to approximately 60.4 % of the overall contaminated soil on-site.

Construction activities during the time frame of this report included: TDF operations and maintenance, wastewater management, feed preparation, debris separation, limited lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling. The contractor also continued to perform off-site and perimeter air monitoring activities.

Excavation continued along the southern end of the excavation face. No drums were encountered during the time frame of this report.

Related site activities included:

* A weekly progress meeting held on Tuesday, October 27, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance. * A junior partnering meeting was held on Wednesday, October 21, 1998 with field management representatives from OHM, the USACE and the PADEP. * A senior partnering meeting and monthly progress meeting was held on Tuesday, October 27, 1998 with representatives from the USEPA, USACE, PADEP and OHM in attendance.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, October 21, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 14, 1998 to 7:00 a.m. Wednesday, Oct. 21, 1998.

The Thermal Destruction Facility (TDF) was on-line for 167.0 hours and processed 6338.9 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.4 % of the time with an average hourly feed rate of 38.0 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 160,225.9 tons of contaminated soil, which amounts to approximately 57.9 % of the overall contaminated soil on-site.

As indicated on last week's report, OHM and the USACE are continuing an investigation of the possible reasons for reduced feed rates to the TDF (i.e. lower than the target value of 40 tons per hour) that had been experienced after the planned maintenance period. The current dual feed screw design and the geophysical properties of the site soil were reviewed as part of the incinerator throughput study. The initial investigation indicated the current addition of lime to prepare the feed soil was contributing to the reduced feed rates (i..e: soils that had a higher quick lime content stuck to the feed screw flights). The soil that stuck to the feed screws caused a greater need for turning power (increased torque) to spin the feed screws, which caused the motor to become overloaded and stop. After several other soil processing methods were investigated, it was decided to process soil with a lower lime content. The processing of lower lime content soils appears to have increased the feed rate to the incinerator to approximately 40 tons per hour. The soil and feed screw investigation will continue throughout the remainder of the project in order to assess the frequently changing soil types existing at the site.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling. The contractor also continued to perform off-site and perimeter air monitoring activities.

Excavation continued along the southern end of the excavation face. Two intact drums were unearthed during the time frame of this report. The intact drums remain in the exclusion zone awaiting sampling for chemical analytical tests and to determine chemical compatibility. One rolloff of decontaminated scrap iron was shipped off-site as scrap.

Related site activities included:

* A weekly progress meeting held on Tuesday, October 20, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* A site tour was given to two groups of students from Bucknell University.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, October 14, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 7, 1998 to 7:00 a.m. Wednesday, Oct. 14, 1998.

The Thermal Destruction Facility (TDF) was on-line for 121.6 hours and processed 3490.5 tons of contaminated soil during the time frame of this report. The TDF was on-line 72.3 % of the time with an average hourly feed rate of 28.7 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 154,432 tons of contaminated soil, which amounts to approximately 55.8 % of the overall contaminated soil on-site.

At the beginning of this reporting period, the TDF remained in a cold shut down mode while repairs were made to the internal and external components of the incinerator. During the down time, OHM replaced the dual feed screws, replaced the covers on the Dry Ash Drag (DAD) conveyors, inspected the internal components of the DAD conveyors, replaced the expansion joints on the scrubber system's pipe lines, performed maintenance on the inside of the acid gas scrubber, and calibrated the TDF monitoring equipment. The incinerator began the heat-up process at 2206 hours on Wednesday evening, October 7, 1998. Once the incinerator had reached the specified temperatures, OHM performed TDF interlock checks in order to test the automatic shutdown systems of the incinerator.

Hazardous waste feed to the incinerator resumed at 0257 hours on Friday, October 9, 1998. The resumption of feed to the incinerator had been delayed because of a local power interruption, which occurred just prior to the scheduled restart of hazardous waste feed. The Thermal Relief Valve (TRV) opened as a result of power loss to the Induced Draft (ID) fan. The backup generator on the thermal pad restored power to the TDF and the TRV was quickly closed. Later that evening, the TRV opened again when power was switched from the backup generator to the local public power grid. For both TRV events, feed to the incinerator had not yet resumed and the proper notifications of the TRV opening were made by the contractor.

Once feed to the TDF had resumed, off-line time to the incinerator was attributed to feed screw stoppages as a result of feed soils sticking to the new dual feed screws. Additional off-line was attributed to feed stoppage during local lightning storm events as a precaution to possible power failure, and feed stopped shortly after the end of the maintenance period in order to perform some additional maintenance to the new feed screws. The lower feed rates to the incinerator are primarily attributed to feed materials sticking to the new feed screws. OHM and the USACE are currently investigating the geophysical properties of the feed soil in relation to current feed screw design. The recent rainy weather and seasonal variation in ambient air temperature are also being evaluated with respect to lower soil feed rates.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling. The contractor also continued to perform off-site and perimeter air monitoring activities.

Excavation continued along the southern end of the excavation face. Two intact drums and one drum carcass were unearthed during the time frame of this report. The intact drums were sampled and placed in overpacks awaiting results of the analytical tests to be performed on the drum contents. Two rolloffs of PPE and site debris (non-hazardous material) were shipped to a hazardous waste landfill. One rolloff of cleaned ductile iron pipe was shipped to a scrap metal dealer.

Related site activities included:

* A weekly progress meeting held on Tuesday, October 13, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* A site tour was given to a group of Junior Engineers in Training for the USACE.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, October 7, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 30, 1998 to 7:00 a.m. Wednesday, Oct. 7, 1998.

The TDF was on-line for 134 hours and processed 4602.2 tons of contaminated soil during the time frame of this report. The TDF was on-line 79.8 % of the time with an average hourly feed rate of 35.3 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 149,802.7 tons of contaminated soil, which amounts to approximately 54.3 % of the overall contaminated soil on-site.

During this period, an instrument air line that aids in monitoring the pressure differential across the bag house was repaired.

Feed to the incinerator was shut down at 2300 hours on Monday, September 5, 1998 to facilitate the replacement of a ruptured expansion joint on a 2" caustic water line pipe from the acid gas scrubber system. In order to keep the scrubber tanks from overheating, and to conduct the repair or the expansion joint, a Thermal Relief Valve (TRV) opening was necessary. Prior to the TRV opening, feed to the incinerator had been shut down for approximately 17 minutes. The necessary authorities were notified immediately of the incident. Because a maintenance shutdown of the TDF was scheduled to begin the morning following the incident, the contractor did not resume feed and started the scheduled plant maintenance.

As stated above, the incinerator was placed in cold shut down mode in order to perform several maintenance tasks to the TDF. During the maintenance period the contractor will replace the dual feed screws, the feed screw housing and cooling jacket, inspect the wet ash drag conveyors, replace the covers on the dry ash drag, replace expansion joints on the acid gas scrubber piping and perform preventative maintenance on several other important components that keep the TDF operating efficiently.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling.

Excavation continued along the southern end of the excavation face near former sediment basin 1. No drums were unearthed during the time frame of this report. The used carbon from the feed preparation building air-handling unit that had been stored on-site was transported off-site.

Related site activities included:

* A weekly progress meeting held on Tuesday, October 6, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* A site visit by Ray Pompe, IT Senior Vice President.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling. The TDF is tentatively scheduled to restart operations on Friday, October 9, 1998.

PREPARED BY: Steve Gibson for David Modricker, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.